1. The mean SAT verbal score is 412, with a standard deviation of 90. Use the Empirical Rule to determine what percent of the scores lie between 412 and 592. (Assume the data set has a bell-shaped distribution.)

34%

81.5%

47.5%

68%

None of the above

Answer by stanbon(75887)

The mean SAT verbal score is 412, with a standard deviation of 90. Use the Empirical Rule to determine what percent of the scores lie between 412 and 592. (Assume the data set has a bell-shaped distribution.)
34%
81.5%
47.5%
68%
None of the above
-----
Find how many std deviations 592 is from the mean.
(592-412)/90 = 580/90 = 6.444
-----
Ans: 50%
---------------
Cheers,
Stan H.
---------------